OpenStudy (anonymous):

[3.06] Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

f(x) = 2(x + 6) 22 = 2(x + 6) 22/2 =x+6 11 =x+6 x=11-6 x=______

OpenStudy (anonymous):

5?

OpenStudy (nurali):

Correct

OpenStudy (anonymous):

WHATS A RANGE? @Nurali